OpenStudy (vera_ewing):

PLEASE HELP ALGEBRA 2! What is a solution of x^2 - 7x = -5?

OpenStudy (dtan5457):

Make it a quadratic so x^2-7x+5=0

OpenStudy (dtan5457):

Do you know how to solve these?

OpenStudy (vera_ewing):

Nope. But I have to submit this in 2 minutes so I have to hurry :/

OpenStudy (dtan5457):

well luckily for you this one is unfactorable so you have to complete the square or quadratic formula..

OpenStudy (dtan5457):

let me do it on paper..

OpenStudy (vera_ewing):

Ok...

OpenStudy (dtan5457):

is this really due in 2 minutes ??

OpenStudy (vera_ewing):

Yep. Actually 1 minute now...

OpenStudy (vera_ewing):

Well now the time is up :/

OpenStudy (dtan5457):

Sorry...do you still want the answer to know how to do this?

OpenStudy (vera_ewing):

Yes please.

OpenStudy (dtan5457):

Ok, let me write it out for you

OpenStudy (vera_ewing):

Thanks.

OpenStudy (dtan5457):

x^2-7x+5=0 since it is unfactorable, you must complete the square. 1.subtract c value to both sides (5) you get x^2-7x=-5 To get a new c value that factorable, you use the formula (b/2)^2 -7/2=3.5^2=12.25 add that to both sides and you get: x^2-7x+12.25=7.25 factor the trinomial which would be (x-7/2)^2=7.25 take out the square by making it x-7/2=the square root of 7.25 add the -7/2 to the other side to get

OpenStudy (dtan5457):

\[x= \frac{ 7 }{ 2 }\pm \sqrt{29/4}\]

OpenStudy (dtan5457):

@vera_ewing

OpenStudy (dtan5457):

I don't know if you ever did this before, but it might be a pain to understand (it took forever for me to get )

OpenStudy (vera_ewing):

Ok that all makes sense but what would be the answer?

OpenStudy (dtan5457):

the equation i just posted ^^

OpenStudy (vera_ewing):

But that isn't one of my answer choices...

OpenStudy (dtan5457):

Post your answer choices

OpenStudy (vera_ewing):

Those are all of my answer choices ^

OpenStudy (dtan5457):

It looks like they used the quadratic formula instead of completing the square, let me try with the quadratic formula.

OpenStudy (vera_ewing):

Ok

OpenStudy (dtan5457):

If you use quadratic formula your final answer would be the second link you posted.

OpenStudy (vera_ewing):

Thank you so much. I really appreciate your help :)

OpenStudy (dtan5457):

Both completing the square and quadratic formula work, so both my answers are theoretically correct, but in this case they used the quadratic formula...which is actually a billion times easier in this case. My bad.

OpenStudy (vera_ewing):

That's alright :) thanks for taking the time to do it both ways!

OpenStudy (dtan5457):

No problem, feel free to tag me into any more questions you may have :)